slipknot3066:

Point Y is the midpoint of segment XZ. If XY = 2(3x +1) and YZ = 5x + 22, find the value of x. x = 20 x = 24 x = 18 x =11

slipknot3066:

@Narad

Narad:

\[XY=YZ\] \[2(3x+1)=5x+22\] \[6x+2=5x+22\] \[6x-5x=22-2\] \[x=20\] The answer is x=20